OpenStudy (valiyuh):

what is the equation of the line with the points (-4,8) and (0,0) ? i need help geometry.! Im stuck

OpenStudy (anonymous):

Basically, to find an equation of a line you can use the point-slope formula in which you need one point on the line and the slope. This is given as \[y - y1 = m (x-x1)\] in which (x1, y1) is any point on the line and m is the slope. To find the slope, \[m = (y1 - y2)\div(x1 - x2) \] (x1, y1) and (x2, y2) being points on the line. So in your case, m = (8-0)/(-4-0) = -2. So the equation is y - 0 = -2 (x-0) further simplified to just y = -2x

OpenStudy (valiyuh):

oh okay thank you !